Bank Math
কর্তৃপক্ষ/ফ্যাকাল্টি
জাতীয় নিরাপত্তা গোয়েন্দা(NSI)সহকারী পরিচালক-০৮.১০.২১
Bank Preparation
Probability, Permutation and Combination
জাতীয় নিরাপত্তা গোয়েন্দা(NSI)সহকারী পরিচালক-০৮.১০.২১
A fair six sided dice is rolled. Find the probability of getting an odd number or a number less than 4.
বিকল্পসমূহ (নিজের উত্তর যাচাই করতে ক্লিক করুন):
সঠিক উত্তর
ক (2:3)
বিস্তারিত ব্যাখ্যা ও বিশ্লেষণ:
Given that a single dice is rolled.
Sample space = {1, 2, 3, 4, 5, 6}
Let P(A) be the probability of getting an odd number, where A = {1, 3, 5}
Let P(B) be the probability of getting a number less than 4, where B = {1,2,3}
A ⋂ B ={1, 3}
So, P(A) = 3/6 = 1/2
P(B) = 3/6 = 1/2
P(A ⋂ B) = 2/6 = 1/3
Let P be the required probability of getting an odd number or a number less than 4
P = P(A ⋃ B) = P(A) + P(B) - P(A ⋂ B)
Or, P(A ⋃ B) = 1/2 + 1/2 - 1/3
Or, P(A ⋃ B) = 1 - 1/3
Or, P(A ⋃ B) = 2/3
Therefore, the probability of rolling an odd number or a number less than 4 is 2/3.